Calculate the number of atoms in one molecule of CH3CH2OH
1
Expert's answer
2020-02-17T07:07:53-0500

Number of atoms in one molecule"=" Total number of molecules of every element"\times" Avagadro's Number

"=(1+3+1+2+1+1)\times 6.022\times 10^{23}"

"=54.198\times 10^{23}"